OpenStudy (anonymous):

please help me .. lim x-> infinity (x-squareroot x ssquare minus 3x )

OpenStudy (anonymous):

\[\lim_{x\to\infty}\left(x-\sqrt{x^2-3x}\right)\cdot\frac{x+\sqrt{x^2-3x}}{x+\sqrt{x^2-3x}}\] \[\lim_{x\to\infty}\frac{x^2-(x^2-3x)}{x+\sqrt{x^2-3x}}\] \[\lim_{x\to\infty}\frac{3x}{x+\sqrt{x^2-3x}}\] In the denominator, factor a \(\sqrt{x^2}\), so that \(\sqrt{x^2-3x}=\sqrt{x^2}\sqrt{1-\dfrac{3}{x}}\). \[\lim_{x\to\infty}\frac{3x}{x+\sqrt{x^2}\sqrt{1-\frac{3}{x}}}\] \(\sqrt{x^2}=|x|=x\), since \(x\to\infty\) (\(x\) is positive, so \(|x|=x\)). \[\lim_{x\to\infty}\frac{3x}{x+x\sqrt{1-\frac{3}{x}}}\]

OpenStudy (anonymous):

oh . so the last answer is ?

OpenStudy (anonymous):

thanks a lot ..

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

this is about l'hospital rule ..

OpenStudy (anonymous):

So you have to use it? I don't see why, since you get the answer right away without it.

OpenStudy (anonymous):

this is my assignment frm my lecturer ..

OpenStudy (anonymous):

At this point \[\lim_{x\to\infty}\frac{3x}{x+\sqrt{x^2-3x}}\] you get the indeterminate form \(\dfrac{0}{0}\) so you could apply the rule then. You'll get something that will have to be dealt with in a similar way, where you factor out \(\sqrt{x^2}\) and so on.

OpenStudy (anonymous):

oh , alright , thanks ..